OpenStudy (anonymous):

Which expression or phrase is equivalent to 9

OpenStudy (anonymous):

3 + 3 + 3 = 9

OpenStudy (anonymous):

\[\sqrt{81}=9\]

OpenStudy (anonymous):

I think we are lacking some information here....

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Which expression or phrase is equivalent to 9? 33 3 • 2 three cubed three squared

OpenStudy (anonymous):

Three squared

OpenStudy (anonymous):

"squared" means "times itself", so three times itself is 3 x 3 which is 9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

estimate 78.65 - 9.5. Explain your method.

OpenStudy (anonymous):

Guess. it's what the question is asking you to do.

OpenStudy (anonymous):

whats the answer though

OpenStudy (anonymous):

Estimate: It is about 80-10, which is 70. To refine the estimate, we might say that we started 3/2 less than 80, and took off 1/2 less than 10, so our refined estimate is one more, or 71. The actual (exact) answer is a touch bigger than that.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

3 • (8 + 2) = 3 • (2 + 8) 2 • (5 • 17) + 12 = (2 • 5) • 17 + 12 2 • (8 + 9) = 2 • 8 + 2 • 9 4 • (7 + 0) = 4 • 7